Solve x2p+y2q=(x+y)z


1
Expert's answer
2021-07-07T05:31:35-0400

Question

Solve x^2p +y^2q=(x+y)z

Solution

x2"\frac {dz}{dx}" +y2"\frac {dz}{dy}" =(x+y)z

Characteristic equation will be

"\frac {dx} {x^2}"="\frac {dy} {y^2}" ="\frac {dz} {(x+y)z}"

"\int\frac {dx} {x^2}" ="\int\frac {dy} {y^2}"

"-\frac 1 x" ="-\frac 1 y" +C1

C1="\frac {x-y} {xy}"

y="\frac {x}{Cx +1}"

The characteristics equation can then be written as;

"\frac {dx} {x^2} =\frac {dz}{(x+y)z}" ="\cfrac {dz} {(x + \cfrac {x} {C_1x +1})z}"

"\frac { C_1x^2 +2x}{x^2(C_1x +1)}"dx ="\frac {C_1x +2} {x (C_1x+1)}" ="\frac {dz} z"

"\int\frac {C_1x+2}{x (C_1x+1)}"dx="\int\frac {dz} z"

But

"\frac {C_1x +2}{x (C_1x+1)}=\frac 1 x +\frac {1}{x(C_1x+1)}"

However,

"\frac {1} {x (C_1x+1)}" ="\frac A x" +"\frac {B}{C_1x+1}"

A(C1x+1)+Bx=1

Therefore,

A=1 and B=-C1

"\frac {1}{x (C_1x+1)}" ="\frac 1 x" +"\frac {-C_1}{C_1x+1}"

"\int" "\frac {C_1x+2}{x (C_1x+1)}" ="\int" "\frac 2 x" -"\int" "\frac {C_1}{C_1x+1}" ="\int\frac{dz}{z}"

2ln|x| -C1ln|C1x+1|=ln|z|+ln|C2|

C2="\frac {x^2}{z(C_1x+1)}" ="\frac {x^2}{z(1+\frac {x-y}{y})}" ="\frac {yx^2}{zx}" ="\frac {xy }{z}"

Answer;

F("\frac {x-y}{xy}","\frac {xy} z")=0